One number is 15 time greater than a second number. When the smaller number is subtracted from twice the larger number the difference is 75. Find the smaller number
Answer by jorel555(1290)

Let your smaller number be n; then the larger number is 15n
2(15n)-n=75
29n=75
n=75/29
15n=1125/29!!!!!!!!!!!!